There are balloons for sale at the fair. 4/10 of the balloons are blue. 2/5 are red, and the rest are yellow. If there are 8 yellow balloons, how many balloons are there in all?
Answer by Theo(13342)

4/10 of the balloons are blue.
2/5 are red.
the rest are yellow.

there are 8 yellow balloons.

4/10 are blue and 2/5 are red.

4/10 + 2/5 = 4/10 + 4/10 = 8/10 of the balloons that are either blue or red.

that leaves 2/10 of the balloons that are yellow.

let x = the number of balloons.

since 2/10 of the balloons are yellow, then the number of yellow balloons is 2/10 * x.

since the number of yellow balloons is 8, then 2/10 * x = 8

solve for x to get x = 10/2 * 8 = 40.

the number of balloons is 40.

4/10 are blue = 4/10 * 40 = 16

2/5 are red = 2/5 * 40 = 16

2/10 are yellow = 2/10 * 40 = 8

total number of balloons is 2 * 16 + 8 = 40

your solution is that the number of balloons in all is 40.